# Momentum accelerated power iterations and the restarted Lanczos method

### v1

@@ 3.1 Example 1
Figure 1 shows reference lines for the predicted rates of convergence given in terms of the Chebyshev polynomials in (2.13) alongside the computed residual convergence of the dynamic momentum algorithm 1.3, and the predicted rates of convergence (2.34), alongside the restarted Lanczos($m$) algorithm 1.4, for $m=8,16,32,64$. In each case $v_{0}$ was chosen as a vector of ones. The plots show the dynamic momentum algorithm converging at its predicted rate, and restarted Lanczos($m$) converging at its predicted rate for $m=8,16,32$, and faster than its rate given by the bound (2.34), for $m=64$. The plots also agree with table 5 which predicts the rates of convergence should agree at $m=m_{cr}=39$: in the bottom left plot we see the rates very close at $m=36$ with the dynamic momentum algorithm slightly faster, and in the bottom right with $m=64$, we see the restarted Lanczos($m$) algorithm faster than both the upper bound for its rate and the dynamic momentum method.
@@ 3.2 Example 2
Let $A=\text{diag}(n:-1:n/2)$, for $n$ even. Then $\varepsilon=1/(n-1)$ and $\varepsilon_{L}=1/(3n/2-1)$. Table 2 shows $m=m_{cr}$ such that restarted Lanczos($m$) attains the same approximate convergence rate as the momentum-accelerated power method with the optimal parameter $\beta$.
@@ 4 Preconditioning
To understand how we may use the momentum-based algorithm 1.2 as a preconditioner for restarted Lanczos, we consider how each of the subdominant eigenmodes in the first approximate eigenvector is reduced by the different algorithms. As discussed in subsection 2.2.2, the static momentum algorithm 1.2 with parameter $\beta=\lambda_{\ast}^{2}/4$ decays all subdominant modes, those with $|\lambda|<|\lambda_{\ast}|,$ at the same average rate, but with higher frequency oscillation for eigenvalues of decreasing magnitude. As shown in [APZ24], similar results hold for the dynamic momentum algorithm 1.3.
Figures 2 and 3 show the average decay rate of each eigenmode for example 1 in subsection 3.1 with n = 1024, for the power iteration, the dynamic momentum algorithm 1.3, and restarted Lanczos($m$), with $m=16,32$ and $64$. The average convergence rate for each mode was computed as the slope of the regression line found by Matlab’s polyfit function, fit to the residual convergence over an entire simulation. The smallest modes for the power iteration are not shown in the plots as they decay to zero very quickly after which their convergence rate is computed as NaN. As expected, we observe each mode in the power iteration decaying proportional to $\lambda/\lambda_{1}$, and each mode in the dynamic momentum algorithm 1.3 decaying at the same average rate. In contrast, the restarted Lanczos($m$) algorithm 1.4 displays oscillatory behavior with respect to $\lambda/\lambda_{1}$. We observe this behavior is not strictly periodic, but rather increases in frequency with the ratio $\lambda/\lambda_{1}$. We refer to this behavior as quasi-periodic. Figure 2 shows the restarted Lanczos($m$) algorithm with $m=16$ (left) and $m=32$ (right), and figure 3 shows $m=64$, with a detail of the average decay rate of the larger subdominant modes.
Figure 2: Average reduction of subdominant eigenmodes by eigenvalue for the power, restarted Lanzcos($m$), and dynamic momentum algorithms applied to example 1 in subsection 3.1. Left: $m=16$; right: $m=32$. The average convergence rate for each mode was computed as the slope of the regression line found by Matlab’s polyfit function.
Figure 3: Average reduction of subdominant eigenmodes by eigenvalue for the power, restarted Lanzcos($m$), and dynamic momentum algorithms. Left: $m=64$; right: $m=64$ (detail). The average convergence rate for each mode was computed as the slope of the regression line found by Matlab’s polyfit function.
@@
Figure 4: Reduction of subdominant eigenmodes by matrix-vector multiplies $k$ for the power, restarted Lanzcos($m$), dynamic momentum, and preconditioned restarted-Lanczos($m$) algorithms for example 1 of subsection 3.1 with $n=1024$. Left: mode $j=12$ with $\lambda_{j}/\lambda_{1}=1013/1024\approx 0.9893$; right: mode $j=64$ with $\lambda_{j}/\lambda_{1}=961/1024\approx 0.9385$
@@ 5 Numerical examples
Figure 5: Residual convergence of restarted Lanczos($m$) algorithm 1.4, the dynamic momentum algorithm 1.3, the momentum preconditioned Lanczos($m$) algorithm 4.1 denoted Lan(mPx,A), and the power preconditioned Lanczos($m$) algorithm denoted Lan(pPx,A), for example 2 of subsection 3.2, with $n=2048$ and $m=64$ (left), and $m=128$ (right).
In figure 5, we show residual convergence for restarted Lanczos($m$) algorithm 1.4, the dynamic momentum algorithm 1.3, the momentum preconditioned Lanczos($m$) algorithm 4.1 denoted Lan(mPx,A), and the power preconditioned Lanczos($m$) algorithm denoted Lan(pPx,A), for example 2 of subsection 3.2. Here, $A$ is the $3073\times 3073$ matrix given by $A=\operatorname{diag}(n:-1:-n/2)$, with $n=2048$. The dimension of the Krylov subspace used for restarted Lanczos in the left plot is $m=64$, and the right plot shows $m=128$. In subsection 3.2 the predicted Krylov subspace dimension $m_{c}$ where algorithms 1.3 and 1.4 would converge at the same rate was $m_{c}=145$. As seen in figure 5, this appears to be an overestimate as the convergence rates are close at $m=128$. Of interest, in both cases we see the momentum preconditioned algorithm 4.1 converge in fewer matrix-vector iterations than the other algorithms tested.
@@ 5.1 Test set 1
Our first set of test matrices given in table 3 is from [WuSi00].
@@
Table 3: Test set 1, the matrices from [DH11] used in [WuSi00]. All three matrices in this set are symmetric positive definite (SPD). The values of $\lambda_{1}$ and $\varepsilon$ listed for each matrix are computed from algorithm 1.4 with $m=64$, and agreed across all methods that sufficiently converged.
@@ 5.2 Test set 2
Our second set of test matrices from [DH11] is used to benchmark performance in [DSYG18]. Each of these matrices is symmetric. Table 5 specifies whether each matrix is additionally positive definite (SPD), and whether it is labeled as a Cholesky candidate in [DH11]. The values of $\lambda_{1}$ listed agreed across all methods.
@@
Table 5: Test set 2, the matrices from [DH11] used in [DSYG18]. All matrices in this set are symmetric. The values of $\lambda_{1}$ listed for each matrix agreed across all methods.
@@
mp-Lanczos(16) 316 19 481 185 F($2\!\cdot\!10^{-4}$)
mp-Lanczos(32) 230 36 295 165 425
mp-Lanczos(64) 196 67 200 196 196
@@
Table 6: Results for the test set 1 matrices given in 3. The number of matrix-vector multiplies to (absolute) residual convergence $\|Ax-\lambda x\|<10^{-12}$. Runs that converged to a residual tolerance greater than $10^{-12}$ after 5000 iterations are denoted F, and the approximate tolerance they stabilized to is listed. Algorithm 1.3 is denoted dyn moment, algorithm 1.4 is denoted Lanczos($m$), algorithm 4.1 is denoted mp-Lanczos($m$) and algorithm 4.2 is denoted $p$-Lanczos($m$.)
The results shown in table 6 show the momentum preconditioned Lanczos($m$) algorithm 4.1 converging with the fewest matrix-vector multiplies for each matrix, with the power preconditioned version 4.2 matching that efficiency in one case. Of interest, the larger Krylov subspace sizes $m$ did not provide an advantage in all of these test cases. In particular, for c-65, the fastest convergence was found with $m=16$. For this matrix, each run of restarted Lanczos without preconditioning terminated after a maximum number of iterations having nearly but not quite achieved the desired tolerance of $10^{-12}$. This demonstrates that the preconditioning can allow increased accuracy. Similar results were found for Ga3As3H12, which achieved the best convergence with momentum preconditioned algorithm 4.1 using $m=32$.
@@ 6 Conclusion
In this paper we quantified the comparative efficiency between a momentum accelerated power iteration and the restarted Lanczos algorithm for symmetric (or complex Hermitian) eigenvalue problems. We then investigated the momentum accelerated power iteration as a polynomial filter or preconditioning technique to accelerate restarted Lanczos. As shown in [APZ24], while each subdominant mode in the momentum accelerated power iteration converges at the same average rate, the convergence of modes with smaller eigenvalues is more oscillatory. Here we also illustrate that the average convergence of the restarted Lanczos method is quasi-periodic with respect to eigenvalue magnitude. We found the momentum accelerated power iteration, which can be efficiently implemented using the dynamic strategy of [APZ24], has a faster convergence rate that restarted Lanczos($m$) for $m<m_{c}$, where $m_{c}$ depends on the relative spectral gap $\varepsilon=(\lambda_{1}-\lambda_{2})/\lambda_{2}$, and also $\varepsilon_{L}=(\lambda_{1}-\lambda_{2^{\prime}})/(\lambda_{2^{\prime}}-\lambda_{n^{\prime}})$, where $\lambda_{2}$ is the eigenvalue of secondary magnitude, $\lambda_{2^{\prime}}$ is the second largest signed eigenvalue, and $\lambda_{n^{\prime}}$ is the left-most eigenvalue in the spectrum of $A$, which agrees with $\lambda_{n}$ when $A$ is positive definite, but not when $A$ has negative eigenvalues.
Our numerical results confirmed our theoretical results on estimating which method would be more efficient based on a priori knowledge of $\varepsilon$ and $\varepsilon_{L}$. Our numerical results further confirmed that momentum accelerated power iterations are an effective preconditioner for the restarted Lanczos algorithm, improving both the number of matrix-vector multiplies to convergence, and residual accuracy. This suggests it may be promising to study the analogous technique for classes of nonsymmetric matrices using the Arnoldi method in place of Lanczos, and the generalized momentum methods of [CMP25-d, CMP25-r] in place of the Chebyshev-based momentum methods.

### v2

@@ 3.1 Example 1
Figure 1 shows reference lines for the predicted rates of convergence given in terms of the Chebyshev polynomials in (2.13) alongside the computed residual convergence of the dynamic momentum algorithm 1.3, and the predicted rates of convergence (2.34), alongside the restarted Lanczos($m$) algorithm 1.4, for $m=8,16,32,64$. In each case $v_{0}$ was chosen as a vector of ones. The plots show the dynamic momentum algorithm converging at its predicted rate, and restarted Lanczos($m$) converging at its predicted rate for $m=8,16,32$, and faster than its rate given by the bound (2.34), for $m=64$. The plots also agree with table 5 which predicts the rates of convergence should agree at $m=m_{cr}=39$: in the bottom left plot we see the rates very close at $m=32$ with the dynamic momentum algorithm slightly faster, and in the bottom right with $m=64$, we see the restarted Lanczos($m$) algorithm faster than both the upper bound for its rate and the dynamic momentum method.
@@ 3.2 Example 2
Let $A=\text{diag}(n:-1:-n/2)$, for $n$ even. Then $\varepsilon=1/(n-1)$ and $\varepsilon_{L}=1/(3n/2-1)$. Table 2 shows $m=m_{cr}$ such that restarted Lanczos($m$) attains the same approximate convergence rate as the momentum-accelerated power method with the optimal parameter $\beta$.
@@
Table 2: The approximate value of $m$ at which (3.1) is satisfied, meaning restarted Lanczos($m$) and power-momentum converge at approximately the same rate for example 2 in subsection 3.2. In the right-hand column where (3.2) is solved by the Matlab command fzero, NN is taken to be 199.
@@ 4 Preconditioning
To understand how we may use the momentum-based algorithm 1.2 as a preconditioner for restarted Lanczos, we consider how each of the subdominant eigenmodes in the first approximate eigenvector is reduced by the different algorithms. As discussed in subsection 2.2.2, the static momentum algorithm 1.2 with parameter $\beta=\lambda_{\ast}^{2}/4$ decays all subdominant modes, those with $|\lambda|<|\lambda_{\ast}|,$ at the same average rate, but with higher frequency oscillation for eigenvalues of decreasing magnitude. As shown in [2], similar results hold for the dynamic momentum algorithm 1.3.
Figures 2 and 3 show the average decay rate of each eigenmode for example 1 in subsection 3.1 with $n=1024$, for the power iteration, the dynamic momentum algorithm 1.3, and restarted Lanczos($m$), with $m=16,32$ and $64$. The average convergence rate for each mode was computed as the slope of the regression line found by Matlab’s polyfit function, fit to the residual convergence over an entire simulation. The smallest modes for the power iteration are not shown in the plots as they decay to zero very quickly after which their convergence rate is computed as NaN. As expected, we observe each mode in the power iteration decaying proportional to $\lambda/\lambda_{1}$, and each mode in the dynamic momentum algorithm 1.3 decaying at the same average rate. In contrast, the restarted Lanczos($m$) algorithm 1.4 displays oscillatory behavior with respect to $\lambda/\lambda_{1}$. We observe this behavior is not strictly periodic, but rather the frequency depends on the ratio $\lambda/\lambda_{1}$. We refer to this behavior as quasi-periodic. Figure 2 shows the restarted Lanczos($m$) algorithm with $m=16$ (left) and $m=32$ (right), and figure 3 shows $m=64$, with a detail of the average decay rate of the larger subdominant modes.
Figure 2: Average reduction of subdominant eigenmodes by eigenvalue for the power, restarted Lanzcos($m$), and dynamic momentum algorithms applied to example 1 in subsection 3.1 with $n=1024$. Left: $m=16$; right: $m=32$. The average convergence rate for each mode was computed as the slope of the regression line found by Matlab’s polyfit function.
Figure 3: Average reduction of subdominant eigenmodes by eigenvalue for the power, restarted Lanzcos($m$), and dynamic momentum algorithms applied to example 1 in subsection 3.1 with $n=1024$. Left: $m=64$; right: $m=64$ (detail). The average convergence rate for each mode was computed as the slope of the regression line found by Matlab’s polyfit function.
@@
Figure 4: Reduction of subdominant eigenmodes by matrix-vector multiplies $k$ for the power, restarted Lanzcos($m$), dynamic momentum, and preconditioned restarted-Lanczos($m$) algorithms for example 1 of subsection 3.1 with $n=1024$ and $m=64$. Left: mode $j=12$ with $\lambda_{j}/\lambda_{1}=1013/1024\approx 0.9893$; right: mode $j=64$ with $\lambda_{j}/\lambda_{1}=961/1024\approx 0.9385$
@@ 5 Numerical examples
Figure 5: Residual convergence of restarted Lanczos($m$) algorithm 1.4, the dynamic momentum algorithm 1.3, the momentum preconditioned Lanczos($m$) algorithm 4.1 denoted Lan(mPx,A), and the power preconditioned Lanczos($m$) algorithm denoted Lan(pPx,A), for example 2 of subsection 3.2, with $n=2048$ and $m=32$ (left), and $m=145$ (right).
In figure 5, we show residual convergence for restarted Lanczos($m$) algorithm 1.4, the dynamic momentum algorithm 1.3, the momentum preconditioned Lanczos($m$) algorithm 4.1 denoted Lan(mPx,A), and the power preconditioned Lanczos($m$) algorithm denoted Lan(pPx,A), for example 2 of subsection 3.2. Here, $A$ is the $3073\times 3073$ matrix given by $A=\operatorname{diag}(n:-1:-n/2)$, with $n=2048$. The dimension of the Krylov subspace used for restarted Lanczos in the left plot is $m=32$, and the right plot shows $m=145$. In subsection 3.2 the predicted Krylov subspace dimension $m_{c}$ where algorithms 1.3 and 1.4 would converge at the same rate was $m_{c}=145$. As seen in figure 5, this appears to be an overestimate as the convergence plot shows an observably steeper slope for restarted Lanczos($m$) than for dynamic momentum with $m=145$. It is further interesting to notice in this plot that the preconditioning stage shows a steeper slope using the momentum method than the power method, however the resulting decrease in preconditioned Lanzcos is more pronounced with the power method preconditioning with $m=145$. However, in the left plot of figure 5 where $m=32$ which is well within the regime where the dynamic method converges faster, we see the momentum preconditioning gives an overall faster convergence then the power preconditioning.
@@ 5.1 Test set 1
Our first set of test matrices given in table 3 is from [22].
@@
Table 3: Test set 1, the matrices from [6] used in [22]. All three matrices in this set are symmetric positive definite (SPD). The values of $\lambda_{1}$ and $\varepsilon$ listed for each matrix are computed from algorithm 1.4 with $m=64$, and agreed across all methods that sufficiently converged.
@@ 5.2 Test set 2
Our second set of test matrices from [6] is used to benchmark performance in [7]. Each of these matrices is symmetric. Table 5 specifies whether each matrix is additionally positive definite (SPD), and whether it is labeled as a Cholesky candidate in [6]. The values of $\lambda_{1}$ listed agreed across all methods.
@@
Table 5: Test set 2, the matrices from [6] used in [7]. All matrices in this set are symmetric. The values of $\lambda_{1}$ listed for each matrix agreed across all methods.
@@
mp-Lanczos(16) 460 24 650 184 F($2\!\cdot\!10^{-4}$)
mp-Lanczos(32) 322 101 511 165 425
mp-Lanczos(64) 325 83 325 327 196
@@
Table 6: Results for the test set 2 matrices given in 3. The number of matrix-vector multiplies to (absolute) residual convergence $\|Ax-\lambda x\|<10^{-12}$. Runs that converged to a residual tolerance greater than $10^{-12}$ after 5000 iterations are denoted F, and the approximate tolerance they stabilized to is listed. Algorithm 1.3 is denoted dyn moment, algorithm 1.4 is denoted Lanczos($m$), algorithm 4.1 is denoted mp-Lanczos($m$) and algorithm 4.2 is denoted $p$-Lanczos($m$.)
The results shown in table 6 show the momentum preconditioned Lanczos($m$) algorithm 4.1 converging with the fewest matrix-vector multiplies for three of the matrices in the set, with the power preconditioned version 4.2 matching that efficiency in one case and exceeding it in another. Of interest, the larger Krylov subspace sizes $m$ did not provide an advantage in all of these test cases. In particular, for c-65, the fastest convergence was found with $m=16$. For this matrix, each run of restarted Lanczos without preconditioning terminated after a maximum number of iterations having nearly but not quite achieved the desired tolerance of $10^{-12}$. This demonstrates that the preconditioning can allow increased accuracy. Similar results were found for Ga3As3H12, which achieved the best convergence with momentum preconditioned algorithm 4.1 using $m=32$. For Si5H12, the fastest convergence was found with $m=32$, using momentum preconditioning although the iteration count differs by only three between $m=32$ and $m=64$, and similar to power preconditioning with $m=64$. One of the interesting observations from this data set is that the momentum preconditioning can be advantageous for restarted Lanczos for smaller Krylov subspace sizes, achieving the same or better convergence than either restarted Lanczos($m$) alone or with power preconditioning with larger Krylov subspace sizes. Overall this points to efficiency advantages in both iteration count and memory.
@@ 6 Conclusion
In this paper we quantified the comparative efficiency between a momentum accelerated power iteration and the restarted Lanczos algorithm for symmetric (or complex Hermitian) eigenvalue problems. We then investigated the momentum accelerated power iteration as a polynomial filter or preconditioning technique to accelerate restarted Lanczos. As shown in [2], while each subdominant mode in the momentum accelerated power iteration converges at the same average rate, the convergence of modes with smaller eigenvalues is more oscillatory. Here we also illustrate that the average convergence of the restarted Lanczos method is quasi-periodic with respect to eigenvalue magnitude. We found the momentum accelerated power iteration, which can be efficiently implemented using the dynamic strategy of [2], has a faster convergence rate that restarted Lanczos($m$) for $m<m_{c}$, where $m_{c}$ depends on the relative spectral gap $\varepsilon=(\lambda_{1}-\lambda_{2})/\lambda_{2}$, and also $\varepsilon_{L}=(\lambda_{1}-\lambda_{2^{\prime}})/(\lambda_{2^{\prime}}-\lambda_{n^{\prime}})$, where $\lambda_{2}$ is the eigenvalue of secondary magnitude, $\lambda_{2^{\prime}}$ is the second largest signed eigenvalue, and $\lambda_{n^{\prime}}$ is the left-most eigenvalue in the spectrum of $A$, which agrees with $\lambda_{n}$ when $A$ is positive definite, but not when $A$ has negative eigenvalues.
Our numerical results confirmed our theoretical results on estimating which method would be more efficient based on a priori knowledge of $\varepsilon$ and $\varepsilon_{L}$. Our numerical results further confirmed that momentum accelerated power iterations are an effective preconditioner for the restarted Lanczos algorithm, improving both the number of matrix-vector multiplies to convergence, and residual accuracy. This suggests it may be promising to study the analogous technique for classes of nonsymmetric matrices using the Arnoldi method in place of Lanczos, and the generalized momentum methods of [4, 5] in place of the Chebyshev-based momentum methods.